40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 25 15:56:11 2016, gautam, Update, General, PRFPMI locked, arms loss improved PRFPMIlock_25Oct2016.pdfITMYwoes.png
    Reply  Tue Oct 25 18:07:28 2016, Koji, Update, General, PRFPMI locked, arms loss improved 
       Reply  Thu Oct 27 12:06:39 2016, gautam, Update, General, PRFPMI locked, arms loss improved PRG.pdf
          Reply  Thu Oct 27 13:48:20 2016, Koji, Update, General, PRFPMI locked, arms loss improved 
             Reply  Thu Oct 27 23:29:47 2016, ericq, Update, General, PRFPMI locked, arms loss improved TRY_REFL_2d.pngAS.pngREFL.pngseis_sub.pdf
                Reply  Fri Oct 28 01:44:48 2016, gautam, Update, General, PRFPMI model vs data studies 6x
                   Reply  Fri Oct 28 19:13:57 2016, rana, Update, General, PR gain 
Message ID: 12579     Entry time: Tue Oct 25 15:56:11 2016     Reply to this: 12580
Author: gautam 
Type: Update 
Category: General 
Subject: PRFPMI locked, arms loss improved 

[ericq,gautam]

Given that most of the post vent recovery tasks were done, and that the ALS noise performance looked good enough to try locking, we decided to try PRFPMI locking again last night. Here are the details:


PRM alignment, PRMI locking

  • We started by trying to find the REFL beam on the camera, the alignment biases for the 'correct' PRM alignment has changed after the vent
  • After aligning, the Oplev was way off center so that was fixed. We also had to re-center the ITMX oplev after a few failed locking attempts
  • The REFL beam was centered on all the RFPDs on the ASDC table

Post the most recent vent, where we bypass the OMC altogether, we have a lot more light now at the AS port. It has not yet been quantified how much more, but from the changes that had to be made to the loop gain for a stable loop, we estimate we have 2-3 times more power at the AS port now.


PRFPMI locking

  • We spent a while unsuccessfully trying to get the PRMI locked and reduce the carm offset on ALS control to bring the arms into the 'buzzing' state - the reason was that we forgot that it was established a couple of weeks ago that REFL165 had better MICH SNR. Once this change was made, we were readily able to reduce the carm offset to 0
  • Then we spent a few attempts trying to do blend in RF control - as mentioned in the above referenced elog, the point of failure always was trying to turn on the integrator in the CARM B path. We felt that the appearance of the CARM B IN1 signal on dataviewer was not what we are used to seeing but were unable to figure out why (as it turns out, we were locking CARM on POY11 and not REFL11 indecision, more on this later)
  • Eric found that switching the sign of the CARM B gain was the solution - we spent some time puzzling over why this should have changed, and hypothesized that perhaps we are now overcoupled, but it is more likely that this was because of the error signal mix up mentioned above...
  • We also found the DC coupling of the ITM Oplev loops to be not so reliable - perhaps this has to do with the wonky ITMY UL OSEM, more on this later. We usually turn the DC coupling on after dither aligning the arms, and in the past, it has been helpful. But we had more success last night with the DC coupling turned off rather than on.
  • Once the sign flip was figured out, we were repeatedly able to achieve locks with CARM partially on RF - we got through about 3 or 4, each was stable for just tens of seconds though. Also, we only progressed to RF on CARM on 1 attempt, the lock lasted for just a few seconds
  • Unfortunately, the mode cleaner decided to act up just about after we figured all this out, and it was pushing 4am so we decided to give up for the night.
  • The arm transmissions hit 300! We had run the transmission normalization scripts just before starting the lock so this number should be reliable (compare to ~130 in October last year). The corresponding PRG is about 16.2, which according to my Finesse models suggest we are still undercoupled, but are close to critical coupling (this needs a bit more investigation, supporting plots to follow). => Average arm loss is ~150ppm! So looks like we did some good with the vent, although of course an independent arm loss measurement has to be done...
  • Lockloss plot for one of the locks is Attachment #1

Other remarks:

  • Attachment #2 shows that the ITMY UL coil is glitchy (while the others are not). At some point last night, we turned off this sensor input to the damping servos, but for the actual locks, we turned it back on. I will do a Satellite box swap to see if this is a Sat. Box problem (which I suspect it is, the bad Sat. Boxes are piling up...)
  • Just now, eric was showing me the CM board setup in the LSC rack, because for the next lock attempts, we want to measure the CARM loop - but we found that the input to the CM board was POY and not REFL! This probably explains the sign flip mentioned above. The mix-up has been rectified
  • The MICH dither align doesn't seem to be working too well - possibly due to the fact that we have a lot more ASDC light now, this has to be investigated. But last night, we manually tweaked the BS alignment to make the dark port dark, and it seemed to work okay, although each time we aligned the PRMI on carrier, then went back to put the arms on ALS, and came back to PRMI, we would see some yaw misalignment in the AS beam...
  • I believe the SRM sat. box is still being looked at by Ben so it has not been reinstalled...
  • Eric has put together a configure script for the PRFPMI configuration which I have added to the IFO configure MEDM screen for convenience
  • For some reason, the appropriate whitening gain for POX11 and the XARM loop gain to get the XARM to lock has changed - the appropriate settings now are +30dB and 0.03 respectively. These have not been updated in some scripts, so for example, when the watch script resets the IFO configuration, it doesn't revert to these values. Just something to keep in mind for now...
Attachment 1: PRFPMIlock_25Oct2016.pdf  284 kB  | Hide | Hide all
PRFPMIlock_25Oct2016.pdf
Attachment 2: ITMYwoes.png  45 kB  | Hide | Hide all
ITMYwoes.png
ELOG V3.1.3-